OpenStudy (anonymous):

Study the following functions and determine if they are continuous. if not, state where their discontinuities exist and what type

OpenStudy (anonymous):

\[f(x)= \frac{ 5 }{ x^4-16}\]

OpenStudy (anonymous):

Hint: You can't divide by zero. Is it possible for \(x\) to be such that computing the function would require dividing by zero? If so, what value of \(x\) would cause that?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

actually, 3

OpenStudy (anonymous):

I got a big homework. I just need to understand how you go on about in understanding how it is discontinuous or not

OpenStudy (anonymous):

i get the whole indivisible by zero part

OpenStudy (anonymous):

there would be a hole if the denominator was zero

OpenStudy (anonymous):

Frank, help me out over here haha

OpenStudy (anonymous):

You were right the first time: when \(x = 2\) we have \(x^4 = 2^4 = 16\) and the denominator of the formula would be zero. And if the denominator is zero then the formula can't produce an answer, and the function has no value. Remember the definition of a continuous function: \(f(x)\) is continuous at \(x_0\) if \(lim_{x\rightarrow x_0} = f(x_0)\). Which means, among other things, that the function \(f\) has to have a value at \(x_0\). So is this particular function continuous at \(x = 2\), or not?

OpenStudy (anonymous):

I said 3 so it wouldn't equal to zero

OpenStudy (anonymous):

i'm guessing you can't just randomly plug in a number?

OpenStudy (anonymous):

At \(x = 3\) the function would have a value. However, that's not what the problem is asking. The problem is asking you if there are places where the function is not continuous, which means places where the function would *not* have a value. So, again, is there a place where this function doesn't have a value, and thus a place where it's not continuous?

OpenStudy (anonymous):

yes, at x=2

OpenStudy (anonymous):

therefore, the function is discontinuous.

OpenStudy (anonymous):

Correct. The function is not continuous at \(x = 2\).

OpenStudy (anonymous):

what type of discontinuity would it be?

OpenStudy (anonymous):

Remember the three types of discontinuity: 1. Removable discontinuity: The function doesn't happen to have a value at 2 (in this case), but you could figure out what it should be based on the value of the function where \(x\) is close to 2 on either side. 2. Jump discontinuity: The function jumps in value when you go from one side of 2 to the other side. For example: a function where the value of the function is 0 for all \(x \le 2\) and 1 for all \(x > 2\). 3. Infinite discontinuity: The value goes to positive or negative infinity (keeps getting ever larger or smaller) the closer you get to 2. So which of the three sounds like it would be the case here? hint: If you have a calculator, try computing the function for 2.0001:\[f(2.0001) = \frac{5}{2.0001^4 -16}\]

OpenStudy (anonymous):

1562.3828

OpenStudy (anonymous):

OK, now try it for 2.00001.

OpenStudy (anonymous):

therefore, it would be an infinite discontinuity.

OpenStudy (anonymous):

Correct, infinite discontinuity.

OpenStudy (anonymous):

because if you're going to keep increasing x, then it's positive infinity. give me a problem with one of the two please :D

OpenStudy (anonymous):

i'm learning a lot here.

OpenStudy (anonymous):

Oh, one more thing I forgot to add: Is \(x = 2\) the only place where this function is not continuous? Hint: Is there any other value of \(x\) for which \(x^4=16\)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No, \(1^4 = 1 \cdot 1 \cdot 1 \cdot 1 = 1\). Try again :-) Hint: Try negative values of \(x\).

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Correct. This function is not continuous at 2 and -2. What type of discontinuity is at -2?

OpenStudy (anonymous):

Removal discontinuity.

OpenStudy (anonymous):

Removal discontinuity is if the function isn't defined to have a value at -2, but you can figure out what the value of the function should be based on the values of the function when \(x\) is close to -2 on either side. For example, suppose we define the function \(g(x)\) to have value of \(x + 2\) for \(x < -2\) and a value of \(2x + 4\) for \(x > -2\), but we don't define a value at -2 itself. What number would make sense for a value of the function at \(x=-2\)?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

and -3

OpenStudy (anonymous):

No, not sure where you got those numbers. But think of it this way: If \(g(x)=x+2\) for \(x<−2\) and \(g(x)=2x+4\) for \(x>−2\), what would the value of \(g\) be at \(x=-2.001\)? What about at \(x = 1.999\)?

OpenStudy (anonymous):

Sorry, I meant \(x = -1.999\).

OpenStudy (anonymous):

-oo

OpenStudy (anonymous):

Sorry, are you talking about the original function \(f(x) = \frac{5}{x^4-16}\)? You're right, that function goes infinite when \(x\) is close to -2. So that would be a "limit of infinity" discontinuity, same as it would be for values of \(x\) close to 2..

OpenStudy (anonymous):

i don't really understand the equation you just gave me. How would i go on about to solve it

OpenStudy (anonymous):

However, for the second function I gave you, where \(g(x) = x+2\) for \(x < -2\) and \(g(x) = 2x+4\) for \(x > -2\), when \(x\) is close to -2 the function \(g(x)\) is close to 0. So the value of \(g(x)\) at \(x = -2\) should be exactly 0. If the value of \(g(-2)\) is not 0 then the function has a removable discontinuity.

OpenStudy (anonymous):

And the way I figured that out is simply by plugging -2 into the function formulas: if \(x = -2\) then \[x+2 = -2+2 = 0\]and\[2x+4 = 2(-2) + 4 = -4+4 = 0\]

OpenStudy (anonymous):

ahhh all right, that makes more sense!!

OpenStudy (anonymous):

wow, missed the value

OpenStudy (anonymous):

Incidentally, the function \(g(x)\) looks something like the following:

OpenStudy (anonymous):

this is exceptional teaching. All right, it all adds up to me now.

OpenStudy (anonymous):

Glad to hear it. Now it's bedtime for me. Good night!

OpenStudy (anonymous):

Thank you so much, I appreciate the patience and effort you put in here. people like you change my life! all right, have a great night! thanks again and hopefully be seing you in the future

OpenStudy (anonymous):

Will do...